# Science and Technology

### v1

@@ Abstract
We study zero-shot instance navigation, in which the agent navigates to a specific object without using object annotations for training. Previous object navigation approaches apply the image-goal navigation (ImageNav) task (go to the location of an image) for pretraining, and transfer the agent to achieve object goals using a vision-language model. However, these approaches lead to issues of semantic neglect, where the model fails to learn meaningful semantic alignments. In this paper, we propose a Prioritized Semantic Learning (PSL) method to improve the semantic understanding ability of navigation agents. Specifically, a semantic-enhanced PSL agent is proposed and a prioritized semantic training strategy is introduced to select goal images that exhibit clear semantic supervision and relax the reward function from strict exact view matching. At inference time, a semantic expansion inference scheme is designed to preserve the same granularity level of the goal-semantic as training. Furthermore, for the popular HM3D environment, we present an Instance Navigation (InstanceNav) task that requires going to a specific object instance with detailed descriptions, as opposed to the Object Navigation (ObjectNav) task where the goal is defined merely by the object category. Our PSL agent outperforms the previous state-of-the-art by 66% on zero-shot ObjectNav in terms of success rate and is also superior on the new InstanceNav task. Code will be released at https://anonymous.4open.science/r/PSL/.
@@ 1 Introduction
Visual navigation is a fundamental skill for embodied agents to travel efficiently in complex 3D environments [65]. Among different kinds of visual navigation tasks, Zero-shot Object Navigation (ZSON) is a promising task and paves the way for developing a general embodied agent [2, 54, 49, 12, 60, 13, 27, 47, 24], since it requires zero scene object annotations for training. However, the criteria of the current ZSON task is to evaluate whether the agent can go to the object that matches solely the given category, which remains a considerable gap from real-life applications that require the identification of a specific object. Therefore, we propose to extend ZSON to Zero-shot Instance Navigation (ZSIN) task in the HM3D environment [42, 58], which requires the agent to go to a unique object instance given detailed text descriptions. Both the ZSON and ZSIN tasks are very challenging due to the zero-shot constraint, where the object goal, required by the reward function in standard reinforcement learning frameworks such as PPO [45, 53], can not be trivially defined.
Figure 1: Pilot studies on four navigation agents: Layout-Only (LO) agent, Semantic-Only (SO) agent, ZSON [37] agent, and our PSL agent. We intuitively think that the agent with a fixed semantic observation encoder i.e., CLIP, possesses good semantic perception ability, while a learnable observation encoder delivers great navigation capacity. (a) The agent trained on ImageNav task does not necessarily need to learn the semantic information to obtain a high success rate. (b)-(c) Our PSL agent achieves both strong semantic perception and navigation capacity. [^2] The LO agent does not support InstanceNav task due to the lack of a semantic goal encoder.
To achieve zero-shot visual navigation, pioneering works [1, 37] turn to pre-trained vision-language models such as CLIP [41] for its zero-shot capability and propose to train the agent on Image-goal Navigation (ImageNav) pretext task that requires the agent to go to a randomly sampled goal image [65]. To be specific, using the CLIP vision encoder to obtain a semantic goal from the goal image for training, the ZSON agent [37] can be transferred to ObjectNav by switching the semantic goal into the CLIP text embedding. Although the ZSON agent establishes a solid semantic agent with good navigation capacity, we argue that the above pipeline does not deliver a navigation agent with strong semantic perception ability.
@@
In this paper, we propose a Prioritized Semantic Learning (PSL) method to improve the semantic perception and understanding ability of the navigation agent. Our PSL approach is comprised of three parts: a PSL agent architecture, a prioritized semantic training strategy, and a semantic expansion inference scheme. First, we equip the agent with an additional CLIP vision encoder to encodes the semantics in the observation. Subsequently, a Semantic Perception Module (SPM) is proposed to comprehend the semantic differences between the observation and goal images. Second, we take advantage of the entropy-minimization technique to select goal images with clear semantic supervision. Moreover, we relax the reward function to focus more on semantic correspondence rather than strict geometric matching Third, considering that the proposed SPM learns to comprehend the semantic differences of image embeddings during training, we develop a semantic expansion inference scheme that retrieves image embeddings using text queries. The retrieved image embeddings expand the text embeddings with rich visual priors, facilitating more precise indications. Without bells and whistles, the proposed PSL method achieves state-of-the-art performance on both widely-evaluated ObjectNav and our extended InstanceNav benchmarks. For example, in the ObjectNav task, PSL outperforms the ZSON baseline [37] and the ESC method [63] that uses large-language models by 16.9% and 3.2% success rate, respectively. Furthermore, we verify the proposed PSL agent has a stronger semantic perception ability by visualizing the semantic similarities between the agent’s last observation and the goal images in Fig. 2 (c).
Our main contributions are summarized as follows:
@@ 4.3 Prioritized Semantic Training Strategy
Previous literature has pointed out that a large portion of image goals in the ImageNav task lead to ambiguity (e.g. looking at walls). We provide statistical analysis on the
Figure 4: Category distribution of goal images in the ImageNav task.
distribution of ImageNav goal images to support this speculation. In particular, We select 6 object categories that widely exist in HM3D scenes, as well as two additional categories “wall” and “room” to represent images with ambiguous semantic content. Then we use CLIP to perform zero-shot classification on all goal images by selecting the category with maximum image-text similarity. In Fig. 4, in the original HM3D ImageNav dataset, most of the goal images cannot be categorized as common object categories; instead, they are classified as meaningless indications (i.e., “wall” and “room”). We believe that the unreasonable distribution of goal images aggravates the semantic neglect issue.
@@
where ${\bf p}_{c}=\mathrm{softmax}\left(g({\bf v}_{w},{\bf q}_{c})\right)$ is the class probability w.r.t. $c^{\text{th}}$ object category, $g({\bf a},{\bf b})=\tau\cdot\frac{{\bf a}^{T}{\bf b}}{\lVert{\bf a}\rVert_{2}\lVert{\bf b}\rVert_{2}}$ is the scaled cosine similarity function and $\tau$ is the temperature. As shown in Fig. 4, the goal images with ambiguous semantics is significantly reduced. Additionally, image distribution among all object categories becomes more balanced.
@@ 4.4 Semantic Expansion Inference Scheme
In the inference stage, the agent receives a text description of the target object. Existing methods directly replace the image-goal embedding with the text-goal embedding of the input text description to perform zero-shot evaluation. However, the granularity gap [32] between image and text embeddings could considerably impair the evaluation performance [31]. Since it is non-trivial to obtain text annotations for all goal images (28.8M) during training, we opt to bridge the modality gap by augmenting the text queries with semantic-consistent image embeddings from a pre-constructed support set. We name this method semantic expansion inference.
During training, we maintain a support set ${\mathcal{S}}=\{{\bf s}\in\mathbb{R}^{d}\mid{\bf s}_{i}\neq{\bf s}_{j},g({\bf s}_{i},{\bf s}_{j})<\lambda\}$. The threshold is set $\lambda=0.8$ to store representative image embeddings, resulting in around 0.1M vectors. These embeddings approximate the distribution of goal images, and they encoded rich information from a variety of object instances, regarding their detailed attributes (e.g., color, shape, texture, and material). During testing, given a language description of the target object, we leverage the CLIP text encoder to produce a feature ${\bf z}_{\mathrm{T}}$. Then we query a goal embedding using the retrieval operation:
${\bf z}_{\mathrm{R}}=\sum_{{\bf s}_{i}\in{\mathcal{S}}}\frac{\mathrm{exp}(g({\bf z}_{\mathrm{T}},{\bf s}_{i}))}{\sum_{{\bf s}_{j}\in{\mathcal{S}}}\mathrm{exp}(g({\bf z}_{\mathrm{T}},{\bf s}_{j}))}\ast{\bf s}_{i}.$ (4)
@@ 5 Experiments
In this section, we first compare the proposed PSL agent with previous zero-shot methods in the ObjectNav task that target finding arbitrary object instances of the same category, and then investigate how it performs when instructed to navigate to a specific object in the InstanceNav task. We also conduct extensive ablation experiments to empirically study the effectiveness of different modules in our method.
@@ 5.1 Experimental Setup
During pre-training, we utilize the 7,200,000 episodes generated by ZSON [37] from HM3D [58]. We randomly select 4 goal images for each episode from 10 candidates with minimal entropy in Eqn. 2. For ObjectNav evaluation, we adopt the habitat2022 challenge dataset with 2000 episodes and 6 object categories. For InstanceNav, we first evaluate the agent on IIN dataset [20], an image-goal track with 795 unique instances in 1000 test episodes, and then extend it to the text-goal setting.
@@
We follow the configuration in [37] with an agent height of 0.88m, radius of 0.18m, and a single $640\times 480$ RGB sensor with a ${}^{\circ}$ horizontal field-of-view (HFOV) placed 0.88m from the ground. For InstanceNav, in the image-goal setting, the agent receives a goal image captured in a random camera configuration that does not match the agent’s own. In the text-goal setting, the agent receives text description of both intrinsic and extrinsic attributes (e.g., “Find a double bed and its material is a beige bamboo frame. There are only two objects around the bed: a table and a window."). For ObjectNav, the agent receives the category text as instruction like “Find a bed”.
@@ 5.2 Experiments on the ObjectNav Task
In Table 1, we report the zero-shot evaluation results on the HM3D ObjectNav task. Our PSL agent outperforms the ZSON baseline [37] by $16.9\%$ in SR ($42.4\%$ vs. $25.5\%$), thanks to our training strategy and inference scheme design. More importantly, our method exceeds the LLM-based methods in the first time. Compared to the PixelNav method [6] that learns to navigate to a point deduced by a powerful multi-modal LLM [40], our PSL agent achieves $+4.5\%$ improvement in SR. Furthermore, taking into account mapping-based methods that perform LLM-guided frontier-based exploration [59], we see $+6.0\%$ improvements against L3MVN [61] and $+3.2\%$ against ESC [63]. Our method shows great potential to be applied in the real environment as it does not require accurate GPS coordinates to build a map. Besides, our LLM-free architecture is more efficient during inference.
5.3 Experiments on the InstanceNav Task
The ObjectNav task assigns multiple available destinations for each episode. It is not sure whether the agent can distinguish object instances. In comparison, we extend the zero-shot object navigation task to the zero-shot instance navigation in the widely-used HM3D environment, as illustrated in Fig. 5. In this task, the agent needs to understand the detailed attributes of an object instance and navigate to an exact location.
Dataset Preparation.
Previous works on the instance-level object navigation task focus on discrete graph-based environments [64] and are limited by complicated human annotation process [11]. In this work, instead, we utilize the recent success of the Generative Vision-and-Language Models (GVLMs) [40, 35, 51] to automatically build the open-vocabulary text-goal setting for the InstanceNav task on the popular HM3D environment. Specifically, we randomly select a goal view for each validation episode to render an image of the goal object in the Instance Image Navigation (IIN) dataset [20]. Each episode in the IIN dataset corresponds to a unique goal object instance. To specify each object instance, we follow a previous attempt [38] to separate text descriptions into two aspects: Intrinsic Attributes and Extrinsic Attributes. Intrinsic attributes cover inherent characteristics of the object, such as shape, color, and material. Extrinsic attributes describe the environment surrounding the object, which is used to determine instances with similar intrinsic attributes. We instruct a GVLM model (i.e., CogVLM [51]) to generate both types of attributes according to the instance image with a hand-crafted prompt (See the Appendix for details). The original ground truth trajectories are preserved to construct the additional InstanceNav test set. In total, the test set of the text-goal setting comprises 1,000 episodes featuring 795 unique objects across 36 scenes.
@@ 5.3 Experiments on the InstanceNav Task
To demonstrate our PSL agent’s ability to understand complicated text descriptions and navigate to specified instances, we compare it with state-of-the-art zero-shot object-goal navigation methods under the text-goal setting of the InstanceNav task. In Table 2, we find that most of the existing zero-shot object navigation methods struggle in this task, including both map-based methods CoW [11] ($1.8\%$ in SR) and LLM-based method ESC [63] ($6.5\%$). All these methods perform inferior to the methods that receive semantic goals (i.e., ours and ZSON), indicating the effectiveness of end-to-end approaches that are pre-trained on large-scale unlabeled data in this task. Similar to our approach, we also transfer a well-trained ImageNav method (OVRL [57]) to perform the text-goal evaluation using our semantic expansion inference scheme. We find that though good at the ImageNav task, the OVRL model performs poorly when given retrieved image goals, indicating that the learnable goal encoder and observation encoder failed to encode semantic-level information for the InstanceNav task. In comparison, our PSL further brings a significant improvement over the ZSON baseline ($+5.9\%$ in SR) and achieves state-of-the-art results.
@@
Table 2: Comparison with state-of-the-art methods in the text-goal navigation track of the InstanceNav task. We report the baseline results based on the released code and models. ${}^{\dagger}$e perform evaluation with our proposed semantic expansion inference scheme.
@@
Table 3: Comparison with state-of-the-art methods in the image-goal navigation track of the InstanceNav task. ${}^{\dagger}$e reimplement OVRL based on released pre-trained weight. In this track, the “Supervised” mark means human labels on the object and viewpoints are used.

### v2

@@ Abstract
We study zero-shot instance navigation, in which the agent navigates to a specific object without using object annotations for training. Previous object navigation approaches apply the image-goal navigation (ImageNav) task (go to the location of an image) for pretraining, and transfer the agent to achieve object goals using a vision-language model. However, these approaches lead to issues of semantic neglect, where the model fails to learn meaningful semantic alignments. In this paper, we propose a Prioritized Semantic Learning (PSL) method to improve the semantic understanding ability of navigation agents. Specifically, a semantic-enhanced PSL agent is proposed and a prioritized semantic training strategy is introduced to select goal images that exhibit clear semantic supervision and relax the reward function from strict exact view matching. At inference time, a semantic expansion inference scheme is designed to preserve the same granularity level of the goal-semantic as training. Furthermore, for the popular HM3D environment, we present an Instance Navigation (InstanceNav) task that requires going to a specific object instance with detailed descriptions, as opposed to the Object Navigation (ObjectNav) task where the goal is defined merely by the object category. Our PSL agent outperforms the previous state-of-the-art by 66% on zero-shot ObjectNav in terms of success rate and is also superior on the new InstanceNav task. Code will be released at https://github.com/XinyuSun/PSL-InstanceNav.
@@ 1 Introduction
Visual navigation is a fundamental skill for embodied agents to travel efficiently in complex 3D environments [66]. Among different kinds of visual navigation tasks, Zero-shot Object Navigation (ZSON) is a promising task and paves the way for developing a general embodied agent [2, 55, 50, 12, 61, 13, 28, 48, 25], since it requires zero scene object annotations for training. However, the criteria of the current ZSON task is to evaluate whether the agent can go to the object that matches solely the given category, which remains a considerable gap from real-life applications that require the identification of a specific object. Therefore, we propose to extend ZSON to Zero-shot Instance Navigation (ZSIN) task in the HM3D environment [43, 59], which requires the agent to go to a unique object instance given detailed text descriptions. Both the ZSON and ZSIN tasks are very challenging due to the zero-shot constraint, in which the category information of goal objects is not feasible during training.
Figure 1: Pilot studies on four navigation agents: Semantic-Non-dominant (SN) agent, Semantic-Dominant (SD) agent, ZSON [38] agent, and our PSL agent. (a) The agent trained on ImageNav task does not necessarily need to learn the semantic information to obtain a high success rate. (b)-(c) Our PSL agent achieves both strong semantic perception and navigation capacity. [^2] The SN agent does not support InstanceNav task as it lacks a semantic goal encoder.
To achieve zero-shot visual navigation, pioneering works [1, 38] turn to pre-trained vision-language models such as CLIP [42] for its zero-shot capability and propose to train the agent on Image-goal Navigation (ImageNav) pretext task that requires the agent to go to a randomly sampled goal image [66]. To be specific, using the CLIP vision encoder to obtain a semantic goal from the goal image for training, the ZSON agent [38] can be transferred to ObjectNav by switching the semantic goal into the CLIP text embedding. Although the ZSON agent demonstrates solid navigation capabilities, we argue that it lacks strong semantic perception abilities.
@@
In this paper, we propose a Prioritized Semantic Learning (PSL) method to improve the semantic perception and understanding ability of the navigation agent. Our PSL approach is comprised of three parts: a PSL agent architecture, a prioritized semantic training strategy, and a semantic expansion inference scheme. First, we equip the agent with an additional CLIP vision encoder to encodes the semantics in the observation. Subsequently, a Semantic Perception Module (SPM) is proposed to comprehend the semantic differences between the observation and goal images. Second, we take advantage of the entropy-minimization technique to select goal images with clear semantic supervision. Moreover, we relax the reward function to focus more on semantic correspondence rather than strict geometric matching Third, considering that the proposed SPM learns to comprehend the semantic differences of image embeddings during training, we develop a semantic expansion inference scheme that retrieves image embeddings using text queries. The retrieved image embeddings expand the text embeddings with rich visual priors, facilitating more precise indications. Without bells and whistles, the proposed PSL method achieves state-of-the-art performance on both widely-evaluated ObjectNav and our extended InstanceNav benchmarks. For example, in the ObjectNav task, PSL outperforms the ZSON baseline [38] and the ESC method [64] that uses large-language models by 16.9% and 3.2% success rate, respectively. Furthermore, we verify the proposed PSL agent has a stronger semantic perception ability by visualizing the semantic similarities between the agent’s last observation and the goal images in Fig. 2 (c). Our main contributions are summarized as follows:
@@ 4.3 Prioritized Semantic Training Strategy
Previous literature [21, 19] has pointed out that a large portion of image goals in the ImageNav task lead to ambiguity (e.g. looking at walls). We provide statistical analysis on the distribution of ImageNav goal images to support this speculation (see Fig. II in Appendix). In particular, We select 6 object categories that widely exist in HM3D scenes, as well as two additional categories “wall” and “room” to represent images with ambiguous semantic content. Then we use CLIP to perform zero-shot classification on all goal images by selecting the category with maximum image-text similarity. As presented in previous literature, in the original HM3D ImageNav dataset, most of the goal images cannot be categorized as common object categories; instead, they are classified as meaningless indications (i.e., “wall” and “room”). We believe that the unreasonable distribution of goal images aggravates the semantic neglect issue.
@@
where ${\bf p}_{c}=\mathrm{softmax}\left(g({\bf v}_{w},{\bf q}_{c})\right)$ is the class probability w.r.t. $c^{\text{th}}$ object category, $g({\bf a},{\bf b})=\tau\cdot\frac{{\bf a}^{T}{\bf b}}{\lVert{\bf a}\rVert_{2}\lVert{\bf b}\rVert_{2}}$ is the scaled cosine similarity function and $\tau$ is the temperature. As shown in Fig. II in the Appendix, the goal images with ambiguous semantics are significantly reduced. Additionally, image distribution among all object categories becomes more balanced.
@@ 4.4 Semantic Expansion Inference Scheme
In the inference stage, the agent receives a series of text descriptions of the target object. Existing methods directly replace the image-goal embedding with the text-goal embedding of the input text description to perform zero-shot evaluation. However, the granularity gap [33] between image and text embeddings could considerably impair the evaluation performance [32]. Since it is non-trivial to obtain text annotations for all goal images (28.8M) during training, we opt to enrich the text goal with fine-grained visual features to reduce the gap between image-goal pre-training and text-goal evaluation. We name this method semantic expansion inference.
During training, we maintain a support set ${\mathcal{V}}=\{{\bf v}\in\mathbb{R}^{d}\mid{\bf v}_{i}\neq{\bf v}_{j},g({\bf v}_{i},{\bf v}_{j})<\lambda\}$ without using any category information. As we expect the selected embeddings to be as dissimilar as possible from the ones already in the set, we set a threshold $\lambda=0.8$ to filter representative image embeddings, resulting in around 0.1M vectors. These embeddings approximate the distribution of goal images, and they encode rich information from a variety of object instances, regarding their detailed attributes (e.g., color, shape, texture, and material). During inference, given a language description of the target object, we leverage the CLIP text encoder to produce a feature ${\bf z}_{\mathrm{T}}$. Then we query a goal embedding using the retrieval operation:
${\bf z}_{\mathrm{R}}=\sum_{{\bf v}_{i}\in{\mathcal{V}}}\frac{\mathrm{exp}(g({\bf z}_{\mathrm{T}},{\bf v}_{i}))}{\sum_{{\bf v}_{j}\in{\mathcal{V}}}\mathrm{exp}(g({\bf z}_{\mathrm{T}},{\bf v}_{j}))}\ast{\bf v}_{i}.$ (4)
@@ 5 Experiments
In this section, we first compare the proposed PSL agent with previous zero-shot methods in the ObjectNav task that target finding arbitrary object instances of the same category, and then investigate how it performs when instructed to navigate to a specific object in the InstanceNav task. We also conduct extensive ablation experiments to study the effectiveness of different modules in our method empirically.
@@ 5.1 Experimental Setup
During pre-training, we utilize the 7,200,000 episodes generated by ZSON [38] from HM3D [59]. We randomly select 4 goal images for each episode from 10 candidates with minimal entropy in Eqn. 2. For ObjectNav evaluation, we adopt the habitat2022 challenge dataset with 2000 episodes and 6 object categories. For InstanceNav, we first evaluate the agent on IIN dataset [21], an image-goal track with 795 unique instances in 1000 test episodes, and then extend it to the text-goal setting. The agent is tested on 20 scenes excluded from the 800 training scenes.
@@
We follow the configuration in [38] with an agent height of 0.88m, radius of 0.18m, and a single $640\times 480$ RGB sensor with a 79 horizontal field-of-view (HFOV) placed 0.88m from the ground. For InstanceNav, in the image-goal setting, the agent receives a goal image captured in a random camera configuration that does not match the agent’s own. In the text-goal setting, the agent receives text description of both intrinsic and extrinsic attributes (e.g., “Find a double bed and its material is a beige bamboo frame. There are only two objects around the bed: a table and a window."). For ObjectNav, the agent receives the category text as instruction like “Find a bed”.
@@ 5.2 Experiments on the ObjectNav Task
In Table 1, we report the zero-shot evaluation results on the HM3D ObjectNav task. Our PSL agent outperforms the ZSON baseline [38] by $16.9\%$ in SR ($42.4\%$ vs. $25.5\%$), thanks to our training strategy and inference scheme design. More importantly, our method exceeds the LLM-based methods in the first time. Compared to the PixelNav method [6] that learns to navigate to a point deduced by a powerful multi-modal LLM [41], our PSL agent achieves $+4.5\%$ improvement in SR. Furthermore, taking into account mapping-based methods that perform LLM-guided frontier-based exploration [60], we see $+6.0\%$ improvements against L3MVN [62] and $+3.2\%$ against ESC [64]. Our method shows great potential to be applied in the real environment as it does not require accurate GPS coordinates to build a map. Besides, our LLM-free architecture is more efficient.
@@
5.3 Experiments on the InstanceNav Task
The ObjectNav task assigns multiple available destinations for each episode. It is not sure whether the agent can distinguish object instances. In comparison, we extend the zero-shot object navigation task to the zero-shot instance navigation in the widely-used HM3D environment, as illustrated in Fig. 4. In this task, the agent receives a textual description of the targeted instance or an instance image as the goal. It needs to understand the detailed attributes of an object instance and navigate to an exact location.
Dataset Preparation.
Previous works on the instance-level object navigation task focus on discrete graph-based environments [65] and are limited by complicated human annotation process [11]. In this work, instead, we utilize the recent success of the Generative Vision-and-Language Models (GVLMs) [41, 36, 52] to automatically build the open-vocabulary text-goal setting for the InstanceNav task on the popular HM3D environment. Specifically, we randomly select a goal view for each validation episode to render an image of the goal object in the Instance Image Navigation (IIN) dataset [21]. Each episode in the IIN dataset corresponds to a unique goal object instance. To specify each object instance, we follow a previous attempt [39] to separate text descriptions into two aspects: Intrinsic Attributes and Extrinsic Attributes. Intrinsic attributes cover inherent characteristics of the object, such as shape, color, and material. Extrinsic attributes describe the environment surrounding the object, which is used to determine instances with similar intrinsic attributes. We instruct a GVLM model (i.e., CogVLM [52]) to generate both types of attributes according to the instance image with a hand-crafted prompt (See Sec. A in the Appendix for details). The original ground truth trajectories are preserved to construct the additional InstanceNav test set. In total, the test set of the text-goal setting comprises 1,000 episodes featuring 795 unique objects across 36 scenes.
@@ 5.3 Experiments on the InstanceNav Task
To demonstrate our PSL agent’s ability to understand complicated text descriptions and navigate to specified instances, we compare it with state-of-the-art zero-shot object-goal navigation methods under the text-goal setting of the InstanceNav task. In Table 2, we find that most of the existing zero-shot object navigation methods struggle in this task, including both map-based methods CoW [11] ($1.8\%$ in SR) and LLM-based method ESC [64] ($6.5\%$). All these methods perform inferior to the methods that receive semantic goals (i.e., ours and ZSON), indicating the effectiveness of end-to-end approaches that are pre-trained on large-scale unlabeled data in this task. Similar to our approach, we also transfer a well-trained ImageNav method (OVRL [58]) to perform the text-goal evaluation using our semantic expansion inference scheme. We find that though good at the ImageNav task, the OVRL model performs poorly when given retrieved image goals, indicating that the learnable goal encoder and observation encoder failed to encode semantic-level information for the InstanceNav task. In comparison, our PSL further brings a significant improvement over the ZSON baseline ($+5.9\%$ in SR).
Table 2: Comparison with state-of-the-art methods in the text-goal track of the InstanceNav task. We report the baseline results based on the released code and models. We perform evaluation with our proposed semantic expansion inference scheme.
@@
Table 3: Comparison with state-of-the-art methods in the image-goal track of the InstanceNav task. In this track, the “Supervised” mark means human labels on the objects are used. We re-implement OVRL based on released pre-trained weight.